# Polaritonic Control of Blackbody Infrared Radiative Dissociation

### v2

@@ III   Results and Discussion
Microcavity-assisted blackbody infrared radiative dissociation
@@
Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity. The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the modified photon DOS.
@@
In Fig. 2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$). For small microcavity lengths, i.e., $L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH, a mild but consistent enhancement of BIRD rates is observed. At these lengths, an overall increase in the microcavity electromagnetic DOS over a wide frequency range covering several important transitions is associated with the observed enhancement. The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns. This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
@@
The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig. 2 arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$. As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
@@
Figure 2: Ratio of BIRD rate inside a microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for HF and LiH molecules. Red circles highlight specific lengths further discussed in Fig. 3.
@@
We conducted a sensitivity analysis of the dissociation rate to interpret the origin of the oscillatory behavior in Fig. 2 and identify the most consequential radiative transitions involved in the BIRD of HF and LiH. This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details). For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate. Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$. In both cases, the identified overtones take the vibrational mode to the highest energy-bound state (23 for HF and 29 for LiH), from which dissociation rapidly occurs. We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths. See Supporting Information for a brief discussion of results obtained without overtone transitions.
Figure 3: Microcavity photon DOS normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig. 2. The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green). Panel (a) shows that the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced relative DOS at the highlighted overtones. Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones. Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.
@@
Equipped with the knowledge of the most relevant transitions for the investigated BIRD of HF and LiH via the most energetic bound state, we can readily explain the microscopic origin of the intricate pattern in Fig. 2 by considering how the microcavity DOS changes around the previously mentioned overtones. Fig. 3 presents the ratio of the microcavity DOS to the free space for selected values of $L_{C}$ where BIRD suppression or enhancement was verified in Fig. 2 (red circles). Overtone frequencies are highlighted with dotted vertical lines, color-coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity. In Fig. 3(a) ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig. 2. In contrast, when the microcavity length slightly increases to $L_{C}=3.21$ $\mu$m [Fig. 3(b)] the photon DOS is enhanced at the $16\rightarrow 23$ and $19\rightarrow 23$ transitions frequencies, and we observe a corresponding leap in the BIRD rate in Fig. 2. Similarly, in Fig. 3(c) ($L_{C}=9.4~{}\mu\text{m})$, all of the relevant LiH overtones but $21\rightarrow 29$ are suppressed, leading to a reduced BIRD rate in this microcavity relative to free space, whereas Fig. 3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$. Introducing a finite linewidth to the considered radiative transitions would smooth out the oscillations observed in Fig. 2 and dampen their amplitude. Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few microns.
@@
Note that suppression effects are generally minor in Fig. 2 because the dissociative process has many pathways. Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always negligible as the molecular system explores alternative dissociation pathways. Our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
@@
Polariton-assisted blackbody infrared radiative dissociation
@@
To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes. Fig. 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photon frequency-momentum dispersion relation [Fig. 1 (e)] and the electromagnetic DOS. We explore the implications for diatomic BIRD via the most energetic bound state below.
@@
Figure 4: Ratio of polariton-assisted BIRD rates to free space BIRD rates. In the polaritonic case, the diatomic molecule is embedded in a strongly coupled microcavity with variable host material with frequency $\omega_{M}$. The top panels, (a) and (b), correspond to HF, and the bottom panels, (c) and (d), show results for LiH. In (a) and (c), the collective light-matter interaction strength, $\Omega_{R}$, is fixed at 400 and 200 cm-1, respectively. The horizontal dashed gray and green lines indicate where the dissociation rates equal $k_{0}$ and $k_{c}$, respectively. The insets in (a) and (c) show zoomed-in views around the overtones $17\rightarrow 23$ for HF and $22\rightarrow 29$ for LiH. In (b) and (d), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the dependency on the detuning between the host molecule and the diatomic transition energies.
Fig. 4 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$. These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 by fixing $L_{C}$ at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH respectively. As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime. We observe prominent peaks for HF and LiH where the relative dissociation rate is notably enhanced. Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$. Here, we focus on the case where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH. In contrast, the inset of Figs. 4(a) and (c) show that when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed. This suppression occurs because as $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the overtone transition falls into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig. 1e). However, the BIRD suppression is relatively small since, as discussed earlier, the reaction may still proceed through alternative pathways.
The effect of the collective coupling strength ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig. 4(b) and (d). When $\omega_{M}-\omega_{i\rightarrow j}$ is around 0.01-1 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates. However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively. This plateau appears to be insensitive to the magnitude of the Rabi splitting. It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig. 5.
@@
Figure 5: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=1401$ cm-1. The curves show the effect of different Rabi splitting values ($\Omega_{R}$) on the photon-weighted polariton DOS, which is relevant for radiative processes mediated by polaritonic systems. As $\Omega_{R}$ increases, the frequency range below $\omega_{M}$, where the photon-weighted polariton DOS significantly exceeds that of free space, widens. This trend explains the impact of the collective light-matter interaction strength on the observed polariton-assisted bond infrared dissociation (BIRD) enhancement shown in Fig. 4.
@@
In Fig. 5, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=9.6$ $\mu$m and $\omega_{M}=v_{22\rightarrow 29}(\text{LiH})=1401$ cm-1. The case where $\Omega_{R}=0$ represents the weak coupling limit and is included here to directly compare $D_{P}(\omega)$ with $D_{C}(\omega)$. It can be seen from Fig. 5 that the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap. Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq. 8. The various curves in Fig. 5 illustrate how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ expand as $\Omega_{R}$ increases. As a result, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting.
@@
The singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate. However, our results demonstrate that this singularity does not lead to arbitrarily fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting. Therefore, the BIRD rates reported as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here. In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap. Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder [63]. The introduction of a UV cutoff, loss, and dephasing pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig. 4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc).
Including a finite linewidth for the diatomic vibrational transitions relaxes the strict resonance requirement discussed above ($\omega_{M}\rightarrow\omega_{i\rightarrow j}$), and the overall enhancement will depend on the spectral overlap between $\omega_{i\rightarrow j}$ and $D_{P}(\omega)$. Since the molecular linewidth and cavity imperfections are expected to be limiting factors, it is clear that our reported results represent theoretical upper bonds for polariton-assisted BIRD rates for the reaction mechanism investigated here.
@@
Before ending, we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in LiH and HF. Since these chemical species have significantly different energetics, we anticipate the main trends reported in this work also apply to other diatomic and polyatomic molecules, particularly in cases where dissociation occurs through a specific doorway state.

### v3

@@ III   Results and Discussion
III.1      Microcavity-assisted blackbody infrared radiative dissociation via doorway state
@@
Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity. The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to the free space, depending on which transitions are affected by the modified photon DOS.
@@
In Fig. 2(a), BIRD rates in the weak coupling regime are shown relative to the free space rate at $T=400~{}\text{K}$. For small microcavity lengths $L_{C}<25$ $\mu$m, a mild but consistent enhancement of BIRD rates is observed. At these lengths, an overall increase in the microcavity electromagnetic DOS over a wide frequency range covering several important transitions is associated with the observed enhancement. The ratios of microcavity BIRD rates to free space BIRD rates approach 1 as $L_{C}$ increases beyond a few tens of microns. This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
@@
On top of the main feature described above, we also observe an oscillatory pattern in the BIRD rates relative to free space shown in Fig. 2(a). This pattern arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$. As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
@@
Figure 2: (a) Ratio of BIRD rate inside a nearly empty microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for NaLi. Red circles highlight specific lengths further discussed in (b) and (c). Microcavity photon DOS normalized to the free space DOS for $L_{C}=26.3$ and $26.8$ $\mu$m, in (b) and (c), respectively. The dotted vertical lines indicate the location of the molecular transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green).
@@
We conducted a sensitivity analysis of the dissociation rate to interpret the origin of the oscillatory behavior in Fig. 2(a) and identify the most consequential radiative transitions involved in the BIRD of NaLi. This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is on the dissociation process (see Supporting Information for more details). For NaLi dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 54$, with $i\in\{42,43,48\}$ are the most consequential for the BIRD rate. The identified overtones take the vibrational mode to the highest energy-bound state, from which dissociation rapidly occurs. We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths. See Supporting Information for a brief discussion of results obtained without overtone transitions.
@@
Equipped with knowledge of the most relevant transitions for the investigated BIRD of NaLi via the most energetic bound state, we can explain the microscopic origin of the intricate pattern in Fig. 2(a) considering how the microcavity DOS changes around the previously mentioned overtones. Figs. 2(b) and 2(c) present the ratio of microcavity DOS to the free space for selected values of $L_{C}$ where suppression or enhancement of BIRD was verified in Fig. 2(a) (red circles). Overtone frequencies are highlighted with dotted vertical lines, color coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity. In Fig. 2(b) ($L_{C}=26.3$ $\mu$m), the most relevant NaLi overtones are suppressed, and this leads to the local minimum observed in $k_{c}/k_{0}$ in Fig. 2. In contrast, when the microcavity length slightly increases to $L_{C}=26.8$ $\mu$m [Fig. 2(c)] the photon DOS is enhanced at the $42\rightarrow 54$ transition frequency, and we observe a corresponding leap in the BIRD rate in Fig. 2(a). Introducing a finite linewidth to the considered radiative transitions would smooth out the oscillations observed in Fig. 2(a) and dampen their amplitude. Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few tens of microns.
@@
Note that suppression effects are generally minor in Fig. 2(a) because the dissociative process has many pathways. Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always negligible as the molecular system explores alternative dissociation pathways. Our analysis of BIRD in weak coupling reveals two main results that we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
We conclude this section by noting that BIRD simulations with lossy mirrors (SI Sec. 8) show (i) broadening and suppression of the sharp features in Fig. 2 and (ii) enhanced dissociation rates relative to Fig. 2(a) at short $L_{C}$ due to the presence of evanescent modes with significant amplitude near the imperfect metal interfaces[7, 61]. Overall, lossy mirrors suppress the oscillatory behavior of the relative BIRD rate with microcavity length while preserving the order-of-magnitude effects predicted for perfect mirrors. This confirms that, modulo the fast oscillations in Fig. 2(a), the results presented in this Section are representative of experimentally accessible photon resonators.
@@
III.2      Polariton-assisted blackbody infrared radiative dissociation via doorway state
@@
To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes. Fig. 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photon frequency-momentum dispersion relation [Fig. 1 (e)] and the electromagnetic DOS. Importantly, the reactive NaLi molecule remains weakly coupled to the confined electromagnetic field. We explore the implications for diatomic BIRD via the most energetic bound state in the following.
@@
Figure 3: Ratio of polariton-assisted BIRD rates to free space BIRD rates. In the polaritonic case, the diatomic molecule is embedded in a strongly coupled microcavity with variable host material with frequency $\omega_{M}$. In (a), the collective light-matter interaction strength, $\Omega_{R}$, is fixed at 200 cm-1. The horizontal dashed gray and green lines indicate where the dissociation rates equal $k_{0}$ and $k_{c}$, respectively. The minimum difference between $\omega_{M}$ and $\omega_{i\rightarrow j}$ used here is 1 cm-1 to match panel (b). The inset in (a) shows a zoomed-in view around the overtone $42\rightarrow 54$. In (b), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the variation of the relative rate enhancement with the detuning between the host molecule and the diatomic transition energies.
@@
Figure 4: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=26.8~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=380$ cm-1. The curves show the effect of different Rabi splitting values ($\Omega_{R}$) on the photon-weighted polariton DOS, which is relevant for radiative processes mediated by polaritonic systems. As $\Omega_{R}$ increases, the frequency range below $\omega_{M}$, where the photon-weighted polariton DOS significantly exceeds that of free space, widens. This trend explains the impact of the collective light-matter interaction strength on the observed polariton-assisted bond infrared dissociation (BIRD) enhancement shown in Fig. 3.
Fig. 3 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$ and fixed microcavity length $L_{C}$ at 26.8 $\mu\text{m}$ where we observed enhanced rates. In Fig. 3(a), we observe prominent peaks where the relative rates are notably enhanced. These peaks occur when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to the specific overtone transition frequencies ($\omega_{i\rightarrow j}$). The asymmetry in the relative dissociation rate profile around the observed maxima, shown in more detail in the inset of Fig. 3(a), emerges as a result of the formation of the polariton stopgap region. When $\omega_{M}$ is close but smaller than $\omega_{i\rightarrow j}$, this transition falls within the stopgap region and is completely suppressed as the density of states becomes zero at that frequency. Nonetheless, as discussed for the weak coupling regime, a strong suppression of reaction rates is not observed due to the many alternative pathways the reaction can access.
Focusing on the $42\rightarrow 54$ overtone ($\omega_{42\rightarrow 54}=380~{}\text{cm}^{-1}$) which is the most consequential transition for the reaction, the effect of a variation in $\omega_{M}-\omega_{42\rightarrow 54}$ at various collective coupling strengths ($\Omega_{R}$) on BIRD rates is examined in Fig. 3(b). Increasing $\Omega_{R}$ leads to a significant enhancement of the polariton-assisted BIRD rates, reaching a factor of 40 at the large Rabi frequency $\Omega_{R}=$200 cm-1 and small $\omega_{M}-\omega_{42\rightarrow 54}$. As this frequency difference increases, the enhancement drops: at 3 cm-1, the enhancement is about 12-fold; at 5 cm-1 detuning, around 7-fold; and beyond 7 cm-1, the enhancement persists around 4-fold. This reduction in dissociation rate enhancement with increasing $\omega_{M}-\omega_{42\rightarrow 54}$ follows from the behavior of the photon-weighted polariton DOS near the stopgap as we discuss next.
@@
In Fig. 4, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=26.8$ $\mu$m and $\omega_{M}=\omega_{42\rightarrow 54}=380$ cm-1. The case where $\Omega_{R}=0$, representing the weak coupling limit, is included here for direct comparison with $D_{C}(\omega)$. It can be seen from Fig. 4 that the photon-weighted polariton DOS is only significantly different from the bare microcavity DOS in a frequency range around $\omega_{M}$ defined as $\{\omega_{M}-\alpha(\Omega_{R})<\omega<\omega_{M}+\alpha(\Omega_{R})\}$, where $\alpha(\Omega_{R})$ is a positive frequency that depends on $\Omega_{R}$ and satisfies $D_{P}[\omega_{M}\pm\alpha(\Omega_{R})]\approx D_{C}[\omega_{M}\pm\alpha(\Omega_{R})]$. Outside this range, $D_{P}(\omega)$ retains features of the weak coupling case, such as oscillations in $\omega$, which can be understood from Eq. 9. Fig. 4 further illustrates how increasing $\Omega_{R}$ broadens the stopgap region and expands $\alpha(\Omega_{R})$, leading to a wider frequency range for which $D_{P}(\omega)\gg D_{C}(\omega)$. As a result, depending on the Rabi splitting, strong coupling can induce changes in the rates of multiple vibrational transitions.
The singular behavior of $D_{P}(\omega)$ at $\omega_{M}$ is originated by the implicit assumption of our model that off-resonant photons with arbitrarily large energy mix with the strongly coupled matter system excitations and form polaritons. These interactions, albeit weak, lead to the formation of lower polariton (LP) modes with very small photon contents and frequencies arbitrarily close to $\omega_{M}$. This behavior is reflected in the expression for the photon-weighted polariton DOS in Eq. 12, where the photon content appears in the numerator and the polariton group velocity in the denominator. As the frequency $\omega$ approaches $\omega_{M}$, the group velocity vanishes faster than the photon content, resulting in an arbitrarily large photon DOS. This divergence in the DOS manifests itself as an accumulation of LP states near $\omega_{M}$, driven by the off-resonant interaction between matter and the microcavity radiation field. In other words, as the in-plane wave vector $k$ increases, the LP transition frequencies approach $\omega_{M}$, yet they never become exactly equal in our model due to the presence of the stopgap. Consequently, within any arbitrarily small frequency interval $\omega_{M}-\delta<\omega<\omega_{M},~{}~{}\delta>0$, a large number of LP states exist, leading to a singularity in the photon-weighted polariton DOS at $\omega_{M}$.
The off-resonant light-matter interactions responsible for the singularity of $D_{P}(\omega)$ at $\omega_{M}$ can be eliminated by imposing an effective frequency cutoff for polaritons formed from high-energy microcavity modes by setting a minimum nonvanishing $\delta=\omega_{M}-\omega_{42\rightarrow 54}$. For example, when $\delta=10$ cm-1 photon modes across 6 branches with $\omega_{C}(m,q)=945$ cm-1 contribute to $D_{P}(\omega)$, while when $\delta=1$ cm-1, $D_{P}(\omega)$ contains contributions from 15 microcavity branches with $\omega_{C}(m,q)=2780$ cm-1. Therefore, by choosing a nonvanishing threshold for $\delta$ we also set a maximum energy (cutoff) for radiation states that contribute to the photon-weighted polariton DOS. Fig. 3 shows the removal of an increasing number of polaritons formed from highly off-resonant photons reduces the relative BIRD rate enhancement to an extent that depends on the Rabi frequency and the minimum $\delta$.
@@
In experiments, imperfections in the strongly coupled material system and in the electromagnetic device (SI, Sec. 8) smooth the photon-weighted DOS near the stopgap. Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet (UV) cutoff and material disorder [70]. The introduction of a UV cutoff, cavity loss, and dephasing pathways for the matter polarization field a̱re therefore expected to lead to reduced BIRD enhancements relative to those in Fig. 3 depending in a non-universal way on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc). Still, given the modest impact of leaky mirrors on the order of magnitude of the microcavity effect on the BIRD rates discussed in Sec. III.A and SI Sec. 8, we expect the order of magnitude of polariton-assisted BIRD enhancement to be similarly robust to the introduction of microcavity losses characteristic of experimentally accessible devices.
Note the singularity in the photon-weighted polariton DOS at $\omega_{M}$ enables a particular radiative transition to become much faster than the corresponding free space rate. However, it does not lead to arbitrarily fast reaction rates, since the complex multistep reactive process ensures that other transitions quickly become rate-limiting (see the Supporting Information for a saturation analysis). Therefore, the reported BIRD rates obtained when $\omega_{M}$ approaches selected $\omega_{i\rightarrow j}$ (Fig. 3) provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here.
@@
Before ending we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in other diatomic species (see SI). Since these chemical species have significantly different energetics, we anticipate the main trends reported in this work also apply to other diatomic and polyatomic molecules, particularly in cases where dissociation occurs through a specific doorway state.
III.3      Dissociation via radiative bound-continuum transitions
In the previous sections, we examined BIRD mediated by the most energetic bound-state. Gas-phase dissociation may also be promoted by direct decay from any bound-state $i\leq i_{\text{max}}$ into the dissociative continuum. This scenario is examined in this section, where we generalize our previous treatment to include direct (irreversible) dissociation pathways for each bound-state $i$ with rate $k^{i}_{\text{loss}}$ as described in Methods and Supporting Information.
Fig. 5(a) shows the relative dissociation rate, $k_{C}/k_{0}$, for a diatomic molecule inside a nearly empty microcavity as a function of its length. In contrast to the doorway mechanism discussed in Sec. IIIA, which exhibits pronounced oscillations, the behavior here is noticeably simpler, yet may be understood with the same tools previously employed to rationalize the observed trends in Fig. 2. For example, from the normalized photon DOS at microcavity lengths $L=10.5~{}\mu\text{m}$ and $L=15.5$ $\mu$m presented in Figs. 5(b,c) corresponding to local minimum and maximum in Fig. 5(a), respectively, we observe the ten most influential transitions, identified through sensitivity analysis, lie between 330–420 cm-1. At $L=10.5~{}\mu\text{m}$, these transitions fall within a frequency range where the DOS is suppressed relative to free space, whereas at $L=15.5~{}\mu\text{m}$, they match with a region of enhanced DOS. This variation accounts for the observed changes and local extrema in the microcavity dissociation rates in Fig. 5(a).
Figure 5: (a) Ratio of BIRD rate inside a nearly empty microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for NaLi. Microcavity photon DOS normalized to the free space DOS for $L_{C}=10.5$ and $15.5$ $\mu$m, in (b) and (c), respectively. The dotted vertical lines indicate the location of the molecular transition frequencies corresponding to the most relevant overtones, with color coding to distinguish between suppression (red) and enhancement (green).
Note when the diatomic dissociation occurs exclusively via the doorway state (as in Sec. III.A.), a few overtone transitions transferring population to the highest energy bound state control the reaction rate. However, when dissociation can proceed directly from each bound state, sensitivity analysis shows that numerous $\Delta v=+2$ transitions (e.g., $12\rightarrow 14$, $13\rightarrow 15$, and $14\rightarrow 16$) can have a significant impact in the reaction rate. This feature challenges the assignment of dominant reaction pathways, yet relaxes the conditions for enhancement or minor suppression of BIRD in a weakly coupled microcavity.
Fig. 6(a) presents the relative polariton-mediated BIRD rate, $k_{P}/k_{0}$, as a function of the host material frequency $\omega_{M}$ under strong coupling conditions with a collective light-matter interaction strength of $\Omega_{R}$ = 125 cm-1 . Here, $L_{C}$ is fixed at 5 $\mu$m, and decay contributions from all vibrational states are included (unlike the results in Sec. III.B which considered only decay from the highest bound state). In contrast to the isolated sharp peaks observed when the reaction proceeds via the loosest bound state (Fig. 3), the introduction of multiple direct dissociative channels via bound-continuum transitions leads to a broader distribution of rate enhancements over a wide range of $\omega_{M}$. While resonant features persist in the reaction rate profile corresponding to particular overtone transitions (Fig. 3a), the observed enhancements largely reflect cumulative contributions of decay channels proceeding directly from many intermediate vibrational bound states.
To estimate the effect of broad vibrational linewidths on the relative polariton-assisted BIRD rates, we applied locally estimated scatterplot smoothing (LOESS)[71, 72, 73], a generalized moving average method. The resulting LOESS curve, shown as a solid line in Fig. 6, reveals the underlying polaritonic enhancement of the BIRD rates while smoothing over sharp features that arise from approximating the continuum states of the Morse potential using a dense but discrete set of box states. This averaging is also supported by the presence of physical broadening mechanisms, including thermal rotational population and weak rovibrational interactions, which naturally contribute to infrared linewidths.
The smoothed polariton-assisted BIRD rates reveal that at $\Omega_{R}$ = 125 cm-1, enhancement is possible when the host material frequency falls within the 300–500 cm-1 range. A mild suppression compared to the weak coupling case ($k_{P}/k_{C}<1$) is also possible when the host material frequency falls within the 200–300 cm-1 range. We attribute these features to the modification of bound-continuum transition rates ($k_{\text{loss}}^{i}$) as demonstrated in Fig. 6(b) which shows these for bound vibrational levels near the dissociation threshold. When $\omega_{M}=230~{}\text{cm}^{-1}$, the decay rates for intermediate bound states are slightly suppressed compared to weak coupling. In contrast, when $\omega_{M}=400~{}\text{cm}^{-1}$, decay rates for states $i=43$–47 are enhanced due to the modified photon weighted polariton DOS. Therefore, in Fig. 6(a), the minimum near $\omega_{M}=230~{}\text{cm}^{-1}$ arises from suppressed decay channels due to a locally depleted polariton DOS, while the maximum near $\omega_{M}=400~{}\text{cm}^{-1}$ results from enhanced decay rates promoted by enhanced polariton DOS. When the reaction is allowed to proceed through this mechanism, which is sensible for gas phase systems, maximum enhancements are reduced compared to the BIRD mediated by the highest energy bound state (or any particular specific doorway state). However, demanding resonance conditions are also lifted affording a more distributed enhancement profile that may offer a practical advantage for experimental realizations.
Figure 6: (a) Ratio of polariton-assisted BIRD rates to free-space BIRD rates for $\Omega_{R}=125~{}\text{cm}^{-1}$. The horizontal dashed green line corresponds to the weakly coupled microcavity relative rate $k_{C}/k_{0}$. The solid line was computed using a LOESS moving average[71, 72, 73], which highlights the overall trend while smoothing out sharp resonance features. (b) Decay rates from bound vibrational states ($i=40$–50) to the continuum under weak coupling ($L_{C}=5~{}\mu\text{m}$), and under strong coupling for different $\omega_{M}$ values in the polaritonic environment.